OpenStudy (anonymous):

A certain mountain has elevation of 19,710 feet. in 1913, the glacier on this peak covered 7 acres. By 2005, this glacier had melted to only 1 acre. Assume that this glacier melted at a constant rate each year. Find this yearly rate. b) use the answer from part a to write a linear equation that gives the acreage A of this glacier t years 1913.

OpenStudy (jdoe0001):

hmmm so you have a mountain covered with snow at the top back in 1913, the snow was covering 7 acres at the top of the mountain and then in 2005, it only covered 1 acre so it melted 6 acres since 1913 a) is the yearly rate, that is, 2005-1913 = 94 years so in 94 years it melted 6 acres, so \(\bf \cfrac{years}{acres} \) will be the yearly rate

OpenStudy (anonymous):

2005-1913 i got 92

OpenStudy (jdoe0001):

b) is just the equation off that say it melted "Z amount" in 1year Acres = "Z amount" t

OpenStudy (anonymous):

2005-1913=92 92/6=15.33

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

A=-0.06+6

OpenStudy (jdoe0001):

actually ... that was my mistake ... it should be Acres per Year melted, thus \(\bf \cfrac{acres}{years}\)

OpenStudy (jdoe0001):

which is a very small amount

OpenStudy (anonymous):

6/92=0.06

OpenStudy (jdoe0001):

yes, then again, one can say, that there are 43560 square feet in 1 acre, thus in 6 acres it'll be 261360 square feet, and if use that to get a rational value, we'd end up with \(\bf \cfrac{acres}{years} \implies \cfrac{6}{92} \cdot \cfrac{261360}{92} \implies 2840.86 \frac{ft}{year}\)

OpenStudy (jdoe0001):

\(\bf 2840.86 \frac{ft^2}{year}\)

OpenStudy (jdoe0001):

.... actually, the B) part uses Acres too... darn... well , I gather we'll use 0.0652 then so Acreage melted = 0.0652t